OpenStudy (anonymous):

line 3x+y=12 intersects line x+2y=4 at P. Given points Q(0,3) and R(3,9), then equation of the line passing P perpendicular to QR is?

ganeshie8 (ganeshie8):

where are you stuck at ?

hartnn (hartnn):

Steps : (A)Q(0,3) and R(3,9), find the slope of line QR using slope formula. (B)Find slope of line through P, because it is perpendicular to QR. (product of slopes of perpendicular lines = -1) Hence you get the slope, now you only need a point to fins the equation. (C) Solve 3x+y=12 and x+2y=4 simultaneously. This will give you point of intersection P. now you have both slope and point, (D) find the equation of line using \(y-y_1 = m (x-x_1)\)

OpenStudy (anonymous):

do you mean slope is gradient?

hartnn (hartnn):

generally while doing vector calculus we use the term gradient, else they both are same.

OpenStudy (anonymous):

emmm oky i am calculating nw!!

OpenStudy (anonymous):

will the answer be y=3x-12 ??

hartnn (hartnn):

"(A)Q(0,3) and R(3,9), find the slope of line QR using slope formula. " i am sure you got the slope as 2. Now, the required line is PERPENDICULAR to line QR so the slope of the required line will be \(\Large \dfrac{-1}{2}\) got this ?

OpenStudy (anonymous):

kaha chali jaati hai bar bar! :O problem to khatm karke jaati :P (i couldnt understand what are you telling sorry) ( i know whr i went wrong i gt the answer ready its 2y+x-2=0 )

OpenStudy (anonymous):

am i right @hartnn ?

hartnn (hartnn):

you mean 2y+x-4 =0 ?? and i thought you knew hindi :P

hartnn (hartnn):

because 2y+x-2=0 is incorrect ! what did u get as co-ordinates of P